Suppose 2 trains leave the station at the same time traveling in opposite direction. One train travels 10 mph faster than the other. In 4.5 hours the trains are 405 miles apart find the speed of each train
Answer by mananth(12270)

Train I x mph
Train II x+ 10 mph
..
They are moving away from each other

So add up their speed.

combined speed = x+ 1 x + 10
( 2 x + 10 )
Time = 4.5 hours
Distance = 405 miles
Distance = speed * time
( 2 x 10 )* 4.5 = 405
9.00 x + 45 = 405
9 x = 405 + -45
9 x = 360
/ 9
x= 40 mph
Train I = 40 mph
Train II 40 + 10 = 50 mph
m.ananth@hotmail.ca